# Ordering kinetics in the active Ising model

### v1

@@ V Summary and Discussion
We conclude this paper with a summary and discussion of our results. We study the ordering kinetics of the active Ising model (AIM), a flocking model with a non-conserved scalar order parameter, after it is quenched from a disordered high-temperature gaseous phase to the phase-coexistence region and the polar-ordered liquid phase. We observe the formation of connected domains of negative and positive spins similar to the ferromagnetic Ising model in the zero activity diffusive limit of the AIM. But for self-propelled particles, AIM manifests an extensive number of disconnected small clusters of the negative and positive spins which eventually merge to a single, macroscopic liquid domain Solon and Tailleur (2015). The domain evolution morphology is characterized by the equal-time two-point correlation function and its Fourier transform, the structure factor. The scaling of the correlation function exhibits good data collapse, signifying a self-similar nature of the domain growth and the large-$k$ behavior of the scaled structure factor tail shows the Porod’s decay which signifies the smooth spatial structure of the AIM domains. The growth law we extract for the AIM follows the Lifshitz-Cahn-Allen (LCA) growth law Bray (2002, 1994); Puri (2009) $R(t)\sim t^{1/2}$ of the non-conserved scalar order parameter and is unaffected by the system control parameters such as temperature, density, and particle velocity. Unlike the VM Katyal et al. (2020), in AIM, the density domain aligns over the same length scale as the orientation and we do not observe any activity-induced correction to the growth law of non-conserved scalar order parameter discussed in the context of the active polar fluid Dikshit and Mishra (2023). We further investigate the role of diffusion on the AIM growth kinetics. We observe that due to the horizontal biased hopping (along ±$\pm x$-direction) of the AIM clusters, particle diffusion along the vertical ±$\pm y$-direction is the predominant mechanism through which coarsening happens in the AIM. This establishes diffusion as the principal growth mechanism rather than activity, leading to a $R(t)\sim t^{1/2}$ growth law for the AIM. We further solve the AIM hydrodynamic equations via a finite difference scheme and the extracted coarsening length scale validates the growth exponent $\theta\simeq 1/2$ observed in the microscopic simulation.
@@
An interesting extension of this study would be to explore the phase ordering kinetics in flocking models in the presence of disorder as experimental systems always contain both quenched and mobile impurities. The similarity in the growth law of a passive system and its active counterpart is an interesting result and therefore, further studies on the coarsening dynamics of active systems, such as the active Potts model or active clock model where the growth law of the corresponding passive models are well-known Grest et al. (1988); Chatterjee et al. (2018, 2020b), are required to confirm (or contradict) this theoretical observation.
@@ Appendix A Comparison of the growth law for the density and magnetization field
Figure 11: (Color online) $R(t)$ versus $t$ for density and magnetization field yields the same growth exponent 0.5 for a quench in the coexistence regime, $\rho_{0}=3$. (Inset) A similar growth law is extracted for a quench in the liquid regime, $\rho_{0}=10$. Parameters: $L=400$, $\beta=1.1$ and $\epsilon=1$.
@@
Fig. 11 shows, for the coarsening of the AIM, that the two length scales exhibit the same growth law, $R(t)\sim t^{1/2}$ (although the domain size for the density field is marginally larger than the corresponding magnetization field) for a quench into the coexistence region and into the polar ordered liquid regime (see inset of Fig. 11). Therefore, we can conclude that the $R(t)\sim t^{1/2}$ growth law is reasonably universal in the AIM as it neither depends on the quenching regime nor the local order parameter (be it density or magnetization).

### v2

@@ V Summary and Discussion
We conclude this paper with a summary and discussion of our results. We study the ordering kinetics of the active Ising model (AIM), a flocking model with both conserved (density) and non-conserved (magnetization) scalar order parameters, after quenching from a disordered high-temperature gaseous phase to the phase coexistence region and the polar-ordered liquid phase. We observe the formation of connected domains of negative and positive spins similar to the ferromagnetic Ising model in the non-motile diffusive limit of the AIM. But for self-propelled particles, AIM manifests an extensive number of disconnected small clusters of the negative and positive spins which eventually merge to a single, macroscopic liquid domain Solon and Tailleur (2015). The domain evolution morphology is characterized by the equal-time two-point correlation function and its Fourier transform, the structure factor. The scaling behavior of the correlation function demonstrates a good data collapse, indicating the self-similar nature of domain growth. Additionally, the large-$k$ behavior of the scaled structure factor tail exhibits Porod’s decay, supporting the smooth spatial structure of the AIM domains. Although AIM contains both conserved and non-conserved order parameters, we extract the same growth law for the AIM density and magnetization fields, the Lifshitz-Cahn-Allen (LCA) growth law Bray (2002, 1994); Puri (2009) $R(t)\sim t^{1/2}$ of the non-conserved scalar order parameter. This indicates that activity does not affect the growth law of the AIM. This observation aligns with earlier findings in Model A with an external drive where external influences do not alter the growth law at the linear level Bray et al. (2001). Unlike the VM Katyal et al. (2020), in AIM, the density domain aligns over the same length scale as the orientation. Furthermore, we do not observe any activity-induced correction to the growth law of the non-conserved scalar order parameter, as observed in the context of the active polar fluid with a non-conserved vector order parameter Dikshit and Mishra (2023). Further, we investigate the role of diffusion in the growth kinetics of the AIM. We observe that due to the horizontal biased hopping (along ±$\pm x$-direction) of the AIM clusters, particle diffusion along the vertical ±$\pm y$-direction is the predominant mechanism through which coarsening happens in the AIM. This establishes diffusion as the main growth mechanism rather than activity, leading to a growth law $R(t)\sim t^{1/2}$ for the AIM. We further solve the AIM hydrodynamic equations using a finite difference scheme, and the extracted coarsening length scale validates the growth exponent $\theta\sim 1/2$ observed in the microscopic simulation. As a future exercise, one could also consider exploring an analytical approach Bray et al. (2001) to solve the hydrodynamic equations to perceive a better understanding of the coarsening dynamics.
In this paper, we focus on purely local particle interactions. Instead of considering a purely local interaction between particles, one can also consider the presence of interactions between the nearest neighboring sites and investigate whether this short-range spatial structure of the interactions affects the extracted growth law. Upon coarse-graining, coupling the magnetization on a site to those on its four neighboring sites will only modify the magnetization equation, but not the density equation, which will retain its form. Consequently, we expect the same growth law exponent in the presence of nearest-neighbor interactions.
@@
An interesting extension of this study would be to explore the phase ordering kinetics in flocking models in the presence of disorder, as experimental systems always contain both quenched and mobile impurities. The similarity in the growth law of a passive system and its active counterpart is an interesting result. Therefore, further studies on the coarsening dynamics of active systems, such as the active Potts model or the active clock model, where the growth laws of the corresponding passive models are well known Grest et al. (1988); Chatterjee et al. (2018, 2020b), are required to confirm (or contradict) this theoretical observation.
@@ Appendix A Comparison of the growth law for the density and magnetization field
Figure 11: (Color online) $R(t)$ versus $t$ for density (solid red square) and magnetization (solid green diamond) field yields the same growth exponent 0.5 for a quench in the coexistence regime, $\rho_{0}=3$. (Inset) A similar growth law is extracted for a quench in the liquid regime, $\rho_{0}=10$. Parameters: $L=400$, $\beta=1.1$ and $\epsilon=1$.
@@
Fig. 11 shows, for the coarsening of the AIM, that the two length scales exhibit the same growth law, $R(t)\sim t^{1/2}$ (although the domain size for the density field is marginally larger than the corresponding magnetization field) for a quench into the coexistence region and into the polar ordered liquid regime (see inset of Fig. 11). Therefore, we can conclude that the $R(t)\sim t^{1/2}$ growth law is reasonably universal in the AIM as it neither depends on the quenching regime nor the local order parameter (be it conserved density or non-conserved magnetization).